OpenStudy (anonymous):

What is the distance between points (12, 4) and (18, –1)? Round to the nearest tenth. 7.8 6.1 4.7 9.9

hartnn (hartnn):

Distance between points (x1,y1) and (x2,y2) is \(\huge d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\)

hartnn (hartnn):

can you find 'd' now ?

OpenStudy (anonymous):

yes. thanks.

hartnn (hartnn):

welcome ^_^ if you want to verify your answer, post it...i'll verify for u.

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

i got \[\sqrt{-23}\]

hartnn (hartnn):

how ? that isn't correct..

OpenStudy (anonymous):

IM SO CONFUSED.

hartnn (hartnn):

ok, so we have \(\huge d=\sqrt{(12-18)^2+(4-(-1))^2}\) right? lets go step by step what is 12-18 =.... ?

OpenStudy (anonymous):

-6

hartnn (hartnn):

correct, and \((-6)^2 =\)

OpenStudy (anonymous):

-32

OpenStudy (anonymous):

NO 32.

hartnn (hartnn):

? how ? 6*6 =.... ? its not 32

OpenStudy (anonymous):

I MEANT 36

hartnn (hartnn):

then you are correct!

OpenStudy (anonymous):

sorry

hartnn (hartnn):

now whats 4-(-1) =.. ?

OpenStudy (anonymous):

5

hartnn (hartnn):

good! and 5^2 =.. ?

OpenStudy (anonymous):

25.

OpenStudy (anonymous):

OKAY I SEE WHERE I MESSED UP!!

hartnn (hartnn):

so you have \(\sqrt {36+25}=..?\)

hartnn (hartnn):

good :) so tell me now what u get...

OpenStudy (anonymous):

7.8

hartnn (hartnn):

\(\checkmark \)

OpenStudy (anonymous):

:D THANKS!!

OpenStudy (anonymous):

could you help me with one more?

hartnn (hartnn):

sure! post it.